# RATE: Score Reward Models with Imperfect Rewrites of Rewrites

### v1

@@ Abstract
This paper concerns the evaluation of reward models used in language modeling. A reward model is a function that takes a prompt and a response and assigns a score indicating how ‘good’ that response is for the prompt. A key challenge is that reward models are usually imperfect proxies for actual preferences. For example, we may worry that a model trained to reward helpfulness learns to instead prefer longer responses. In this paper, we develop an evaluation method, RATE (Rewrite-based Attribute Treatment Estimators), that allows us to measure the causal effect of a given attribute of a response (e.g., length) on the reward assigned to that response. The core idea is to use large language models to rewrite responses to produce imperfect counterfactuals, and to adjust for rewriting error by rewriting twice. We show that the RATE estimator is consistent under reasonable assumptions. We demonstrate the effectiveness of RATE on synthetic and real-world data, showing that it can accurately estimate the effect of a given attribute on the reward model. Code is available at https://github.com/toddnief/RATE.
@@ 1 Introduction
In the context of large language models (LLMs), reward models evaluate the quality or appropriateness of model outputs, either by assessing individual responses or comparing multiple alternatives. Such models are useful in a variety of settings, including alignment of large language models, ranking output samples (e.g., to use in a best-of-$n$ sampling procedure), or evaluation of LLM performance.
Ideally, reward models would directly and perfectly measure whatever aspect of the output is important—e.g., we might have a reward for mathematical problem solving based on whether the generated response is correct. However, reward models are commonly learned from training data that imperfectly measures somewhat nebulous attributes. For example, a common task is to train a reward model based on human preferences for which of two responses is more helpful. This results in a challenge where, even with a reward model in hand, we are not certain what it is actually rewarding. For example, we might worry that a model trained to reward helpfulness learns to instead simply prefer longer responses (Shen et al., 2023; Park et al., 2024b; Singhal et al., 2024).
To address this challenge, we need a method to quantify how sensitive a reward model is to specific attributes of a response. A straightforward approach would be to collect a dataset of prompt/response pairs, label each response as having or not having the attribute of interest, and then compare the average reward assigned to responses with and without the attribute. However, this approach has the limitation that it does not account for ‘spurious’ correlations that may exist in the data. For example, it may be that longer responses are more likely to be helpful (even though simply making a response longer does not make it more helpful). Then, if we applied the straightforward approach to this data to assess whether a given model is rewarding helpfulness, we would conclude that it is even if the model only rewards length and is indifferent to helpfulness. If we then used this reward model as a proxy for helpfulness in a downstream alignment task, then the actual effect of alignment would be to make responses longer, without (necessarily) affecting helpfulness.
Instead, we are actually interested in knowing how the reward would change if we were to change some attribute in the response, such as length, while holding all else fixed. This is the causal effect of the attribute on the reward. There is a growing literature on estimating the causal effects of attributes of text (Feder et al., 2022; Grimmer et al., 2022; Jin et al., 2022; Chen & Chu, 2023; Gui & Veitch, 2023).
Generally, these provide methods for estimating the causal effect using observational data, where we cannot intervene on the text directly. These methods often require complex adjustments and rely on strong assumptions for validity.
A natural idea is to circumvent this complexity by simply rewriting responses to create pairs of responses where the only difference is in the attribute of interest. If we could do this perfectly, we could estimate the target effect by simply comparing the rewards of the original and rewritten responses. Of course, rewrites cannot be done perfectly.
The contribution of this paper is to develop and demonstrate a rewrite-based method for estimating the causal effect of an attribute of a response, on the reward assigned to that response:
@@
    We develop a practical method of estimating the causal effect of an attribute of a response on reward using imperfect LLM-based rewrites. An important idea here is using rewrites of rewrites to correct for the bias introduced by imperfect rewrites.
@@
    We show that this method is an unbiased and consistent estimator of the causal effect.
@@
    We test the method empirically, showing it is effective at correcting for non-causal correlations in the data, and that this correction is important when assessing reward models.
@@ 2 Setup
Reward models are typically implemented in two ways:
  1. 1.
    As functions $R(x,y)$ that take a prompt $x$ and a response $y$ as inputs and return a real number indicating the quality of the response for the prompt.
  2. 2.
    As functions $R(x,y_{1},y_{0})$ that take a prompt $x$ and two responses $y_{1}$ and $y_{0}$ as inputs and return a real number describing the relative quality of $y_{1}$ compared to $y_{0}$.
Our results apply to both implementations, but we focus on the first for clarity (see Section 6).
Suppose we have a dataset of prompt-completion pairs $\{(x^{i},y^{ij})\}$, where the $x^{i}$ are prompts and the $y^{ij}$ are completions (also referred to as ‘responses’). We have a reward model $R(x^{i},y^{ij})$ that assigns a scalar reward to each prompt-completion pair. We are interested in understanding how the reward model responds to a certain attribute, represented by the function $W$, within the completions. For each prompt-completion pair, we have a binary label $w^{ij}=W(x^{i},y^{ij})\in\{0,1\}$ indicating whether the completion has the attribute of interest.
For example, $W$ might represent helpfulness, which varies based on the context given by the prompt. A recipe could be helpful for questions about cooking but not for questions about history.
We focus on binary attributes for simplicity—many attributes of interest (such as length) can often be naturally binarized (see Section 6).
@@
Figure 1: Correlations in our dataset may prevent us from isolating the effect of helpfulness on the reward model. For instance, helpful responses may tend to be longer.
@@
$\hat{\tau}_{\text{naive}}=\frac{1}{n_{1}}\sum_{(x^{i},y^{ij}):w^{ij}=1}R(x^{i},y^{ij})-\frac{1}{n_{0}}\sum_{(x^{i},y^{ij}):w^{ik}=0}R(x^{i},y^{ik})$
@@
We may view this as a finite sample estimator for the quantity:
$\mathbb{E}[R(X,Y)\leavevmode\nobreak\ |\leavevmode\nobreak\ W=1]-\mathbb{E}[R(X,Y)\leavevmode\nobreak\ |\leavevmode\nobreak\ W=0],$
where the expectation is taken over the distribution from which our evaluation examples are drawn. The problem here is that, even in the infinite data limit, this quantity does not generally isolate the effect of $W$ on $R$. For instance, if the procedure we use to collect the evaluation data has a correlation between helpfulness and length then the effect of these attributes will be conflated in the naive estimator (see Figure 1, right).
@@
$\text{ATE}=\mathbb{E}[R(X,Y(1))-R(X,Y(0))]$
where $X$ is a random variable for the prompt, and $Y(1)$ and $Y(0)$ are potential outcomes for responses. This quantity is the expected change in reward if we were to change the attribute $W$ from 0 to 1, while keeping all other aspects of the response fixed. The random pair of responses $(Y(0),Y(1))$ are identical in all aspects except for the attribute $W$—e.g., if $W$ is helpfulness then each counterfactual response should have the same writing level, sentiment, topic, etc. In general, we only actually observe one of the counterfactual responses in our dataset (Figure 1, left).
@@
Table 1: GPT-4o qualitatively does well at rewriting IMDB responses to change sentiment from negative (W = 0) to positive (W = 1). The first example was selected for illustrative purposes, the latter two were randomly selected from the dataset.
Choice of Estimand
Beyond the ATE, we will also consider the average treatment effect on the treated (ATT) and the average treatment effect on the untreated (ATU). These are defined as:
ATT $=\mathbb{E}\left[R(X,Y(1))-R(X,Y(0))|W=1\right]$
ATU $=\mathbb{E}\left[R(X,Y(1))-R(X,Y(0))|W=0\right]$
Intuitively, if $W=1$ is a helpful response, the ATT measures the change in reward when we take a helpful response and make it less helpful, and the ATU measures the change in reward when we take an unhelpful response and make it more helpful. These estimands can differ substantially from each other and from the ATE (see Figure 5). There is no reason to expect these quantities to align in general, so some thought should be given to which is most relevant to the question at hand. Indeed, even human preferences are often asymmetric (Kahneman & Tversky, 2013), so we might expect reward model preferences to be as well.
3 RATE: Rewrite-based Attribute Treatment Estimators
Whatever our choice of estimand, we need a method to estimate it. Here, we develop a method, RATE, that uses rewrites to estimate the causal effect of an attribute on a reward model. The core idea is to create pairs of responses where the only difference is in the attribute of interest. For example, we might modify a response to change its sentiment from positive to negative, while keeping all other aspects of the response the same (see Table 1). The goal is for these modified responses to directly approximate the unobserved counterfactual responses.
@@ 3 RATE: Rewrite-based Attribute Treatment Estimators
In practice, we implement rewrites using a large language model (LLM). We begin with a labeled dataset containing ground truth binary variables for attributes such as complexity, sentiment, or helpfulness. We then instruct the LLM to rewrite the responses to the opposite state of the binary variable. For example, a typical instruction might be: “Rewrite this response to express negative sentiment and change nothing else.”
We use $\text{Re}(x^{i},y^{ij},w)$ to denote the rewrite operation, which takes a prompt-response pair $(x^{i},y^{ij})$ and a desired attribute value $w$, returning a modified response $\tilde{y}^{ij}$ such that $W(x^{i},\tilde{y}^{ij})=w$.
@@
For instance, when rewriting for ‘helpfulness’, we might instruct the LLM to “Rewrite this response to be more helpful”, or instruct it to “Rewrite this response to be more helpful, providing additional relevant information or clarification.” In this example, the second instruction makes the meaning of ‘helpful’ more precise. Generally, changing the instruction changes the nature of the rewrites generated, and thus changes the attribute that is being modified.
This is inevitable. Ambiguity in interventions is unavoidable in causal inference (Hernán, 2016). In our context, there is subjectivity in what helpfulness, complexity, or sentiment actually mean. An advantage of the rewrite approach is that it allows us to use natural language to specify, as clearly as possible, what property we are actually trying to modify. We can understand whether our instructions are having the intended effect by qualitatively examining the rewritten outputs and checking that they vary the attribute of interest while leaving the rest of the response unchanged. In practice, finding effective rewrite instructions requires an iterative cycle of generating rewrites, examining the responses, and adjusting the rewrite prompt to be more clear and specific.
Attribute W=1 W=0 Rewrite Prompt
Table 2: Example rewrite prompts from experiments with sentiment and length as the target attribute. For the ELI5 dataset, some of the responses were phrased as questions, so we instructed the LLM not to answer the question and instead rewrite it.
@@
If the rewrites produced perfect counterfactuals, it would then be straightforward to estimate the causal effect of the attributes. Namely, we could compare the rewards of the original responses to the rewards of the rewrites. However, the rewrites are often imperfect, modifying off-target attributes. These off-target modifications may affect the reward, causing the simple comparison to be misleading. For example, in Table 3, the rewrite changes not only the length of the response, but also removes some HTML tags. Changing the off-target attributes can affect the reward, leading to a biased estimate of causal effects.
Original (W = 1) Rewrite (W = 0)
…I really had to see this for myself.<br /><br /> The plot is centered around a young Swedish drama student named Lena… …so I had to see it for myself. The plot centers around Lena, a Swedish drama student …
Table 3: Excerpt from rewriting IMDB responses to change length from long $(W=1)$ to short $(W=0)$. HTML tags (an off-target attribute) are removed in the rewrite.
Mathematically, whenever we rewrite some response $y^{ij}$ (to $W=w$), we introduce some error $\varepsilon^{ij}_{w}$ in the reward because of our inability to perfectly produce the counterfactual $y^{ij}(w)$, which ought to differ from the original response only with respect to the target attribute. Define this error as:
$\varepsilon^{ij}_{w}=R(x^{i},\text{Re}(x^{i},y^{ij},w))-R(x^{i},y^{ij}(w))$
We would like to correct for these errors. Yet the whole point of the rewrites is to approximate the counterfactuals $y^{ij}(w)$, so we cannot directly measure $\varepsilon^{ij}_{w}$.
Original Rewrite Rewrite of Rewrite
W = 0, Reward: 0.14 W = 1, Reward: 0.12 W = 0, Reward: 0.16
W = 0, Reward: 0.13 W = 1, Reward: 0.17 W = 0, Reward: 0.16
W = 1, Reward: 0.135 W = 0, Reward: 0.134 W = 1, Reward: 0.139
Table 4: Whether for a rewrite or a rewrite-of-a-rewrite, GPT-4o uses well-formatted text and a slightly formal tone. Here, W is length; samples are drawn from the ELI5 dataset, scored using ArmoRM, and truncated to 100 characters for display. The first was selected for illustrative purposes, the latter two were randomly selected from the dataset.
@@
Surprisingly, the solution is to introduce more noise. Instead of comparing a rewrite to the original response, we compare it to the rewrite of the rewrite, thereby canceling out off-target changes introduced by the rewrite process. That is, rather than selecting (original, rewrite):
$\tilde{\tau}^{ij}=\begin{cases}R(x^{i},y^{ij})-R(x^{i},\text{Re}(x^{i},y^{ij},1)),&\text{if }w^{ij}=1\\ R(x^{i},\text{Re}(x^{i},y^{ij},0))-R(x^{i},y^{ij}),&\text{if }w^{ij}=0\end{cases}$
@@
$\hat{\tau}^{ij}=\begin{cases}R(x^{i},\text{Re}(\text{Re}(x^{i},y^{ij},0),1))-R(x^{i},\text{Re}(x^{i},y^{ij},0)),&\text{if }w^{ij}=1\\ R(x^{i},\text{Re}(x^{i},y^{ij},1))-R(x^{i},\text{Re}(\text{Re}(x^{i},y^{ij},1),0)),&\text{if }w^{ij}=0\end{cases}$
The motivation is that the off-target changes introduced by the rewrite process will, in expectation, cancel out when we are comparing two things in ‘rewrite space’. For example, the tendency for LLMs to produce well-formatted text will affect both the first rewrite and the rewrite of the rewrite (as shown in Table 4), so the overall contribution of this off-target change will cancel out. This approach yields the Rewrite-based Attribute Treatment Estimators (RATE) for the ATT, ATU, and ATE:
@@
1:Input: Dataset $\{(x^{i},y^{ij},w^{ij})\}$, reward model $R$, function $\text{Re}()$
2:Return: Estimates $\hat{\tau}_{\text{ATT}}$, $\hat{\tau}_{\text{ATU}}$, $\hat{\tau}_{\text{ATE}}$
3:Initialize $n_{1}\leftarrow\sum_{i,j}\mathbb{I}[w^{ij}=1]$, $n_{0}\leftarrow\sum_{i,j}\mathbb{I}[w^{ij}=0]$
4:$\hat{\tau}_{\text{ATT}}\leftarrow\frac{1}{n_{1}}\sum\limits_{(i,j):w^{ij}=1}[R(x^{i},\text{Re}(\text{Re}(x^{i},y^{ij},0),1))-R(x^{i},\text{Re}(x^{i},y^{ij},0))]$
5:$\hat{\tau}_{\text{ATU}}\leftarrow\frac{1}{n_{0}}\sum\limits_{(i,j):w^{ij}=0}[R(x^{i},\text{Re}(x^{i},y^{ij},1))-R(x^{i},\text{Re}(\text{Re}(x^{i},y^{ij},1),0))]$
6:$\hat{\tau}_{\text{ATE}}\leftarrow\frac{n_{1}}{n_{0}+n_{1}}\hat{\tau}_{\text{ATT}}+\frac{n_{0}}{n_{0}+n_{1}}\hat{\tau}_{\text{ATU}}$
7:return $\hat{\tau}_{\text{ATT}}$, $\hat{\tau}_{\text{ATU}}$, $\hat{\tau}_{\text{ATE}}$
In practice, we may not have $w^{ij}$ for all examples, so we can use a classifier to predict $w^{ij}$ from $x^{i}$ and $y^{ij}$, and then use the classifier’s predictions in the RATE estimators.
4 Theoretical Analysis of RATE
Under reasonable assumptions, RATE is a consistent estimator of the average treatment effect.
Latent Variable Model
To analyze the rewrite operation, we need to conceptualize how different aspects of a response might change during rewriting. Imagine a response as having three types of attributes: the target attribute we want to change (like sentiment), attributes that should remain constant (like topic), and attributes that might unintentionally change (like specific wording). We can formalize this idea using a latent variable model:
$Y=Y(W,Z,\xi)$
where:
  • $Y$ is the observed response
  • $W$ is the target attribute we aim to manipulate (e.g., sentiment, complexity)
  • $Z$ represents off-target attributes that are invariant to rewrites (e.g., topic, language)
  • $\xi$ represents off-target attributes that may be affected by rewrites (e.g. grammatical structure)
Intuitively, we expect some off-target attributes $Z$ to remain unchanged during rewrites. For example, if we ask a large language model to change the sentiment of an English text, we don’t expect it to suddenly produce Korean. However, other off-target attributes $\xi$ may change: for instance, grammar and punctuation might be corrected.
Unbiasedness and Consistency of RATE
To establish that RATE is a sound estimator of the causal effect we need some additional assumptions:
  1. 1.
    We assume that the reward model can be decomposed additively:

### v2

# RATE: Causal Explainability of Reward Models with Imperfect Counterfactuals

@@ Abstract
Reward models are widely used as proxies for human preferences when aligning or evaluating LLMs. However, reward models are black boxes, and it is often unclear what, exactly, they are actually rewarding. In this paper we develop Rewrite-based Attribute Treatment Estimator (RATE) as an effective method for measuring the sensitivity of a reward model to high-level attributes of responses, such as sentiment, helpfulness, or complexity. Importantly, RATE measures the causal effect of an attribute on the reward. RATE uses LLMs to rewrite responses to produce imperfect counterfactual examples that can be used to measure causal effects. A key challenge is that these rewrites are imperfect in a manner that can induce substantial bias in the estimated sensitivity of the reward model to the attribute. The core idea of RATE is to adjust for this imperfect-rewrite effect by rewriting twice. We establish the validity of the RATE procedure and show empirically that it is an effective estimator. Code is available at https://github.com/toddnief/RATE.
Machine Learning, ICML
@@ 1 Introduction
Reward models (RMs) play a critical role in aligning large language models (LLMs) with desired behaviors. These models evaluate the quality of LLM outputs and are widely used in, e.g., post-training, inference-time sampling adjustment (through best-of-$n$ sampling), and LLM evaluation. However, reward models, typically implemented as fine-tuned LLMs, are black boxes. Accordingly, it can be difficult to understand what a reward model is actually rewarding. Despite their importance, explainability research for RMs remains relatively underexplored (Lambert et al., 2024).
To understand reward models we would like to be able to measure their responsiveness to high-level attributes of responses, such as helpfulness, correctness, or sentiment. Such responsiveness measurements would serve as an important diagnostic tool. The purpose of the present paper is to develop such a measurement procedure.
Naively, one might attempt to estimate RM responsiveness to an attribute by comparing the average reward assigned to responses with and without the attribute using a labeled dataset. However, this approach is flawed: it conflates the target attribute’s influence with correlations present in the evaluation data. For instance, suppose we are measuring how much the RM responds to “sentiment” and, unbeknownst to us, highly-negative samples in our evaluation data tend to have more typos and sloppy formatting. Then if we naively average the reward over the negative samples, and again over the positive samples, we’ll inadvertently measure the effect of not just sentiment, but typos and formatting as well. In particular, this means we can measure a large effect of an attribute on the RM, even if the RM is completely insensitive to that attribute.
For measurements of attribute influence to be meaningful, they must be isolated from confounding factors and spurious correlations in the evaluation data. To address this, we propose formalizing RM explainability as the causal effects of attributes on the reward. Hence, the goal is to determine how a reward would change if we could modify a response to alter only the attribute of interest while holding all other factors constant. This counterfactual perspective isolates the causal relationship, making it possible to disentangle an attribute’s true effect from confounding factors.
There is a growing literature on estimating the causal effects of attributes of text (Feder et al., 2022; Grimmer et al., 2022; Jin et al., 2022; Chen & Chu, 2023; Gui & Veitch, 2023). Generally, this work provide methods for estimating the causal effect by either making strong assumptions on the evaluation data or compensating for potential confounding factors.
Notice that if we had access to counterfactual pairs of responses (i.e., pairs where the only difference is the attribute of interest), we could estimate the target effect by simply comparing the rewards assigned to each response. A natural idea is to use LLMs to generate such pairs by rewriting responses to change the target attribute. If the rewrites were perfect, we could directly measure the causal effect of the attribute on the reward. However, in practice, LLMs produce imperfect rewrites, changing off-target attributes as well. These imperfections can substantially bias the estimated causal effect, as we will see in Section 5.
The contribution of this paper is to develop a procedure that allows us to use the imperfect rewrites from LLMs to get a correct estimate of the causal effect of an attribute on a reward model. The development is as follows:
@@
    We formalize RM explainability as an Average Treatment Effect (ATE) of high-level attributes of responses, on the reward assigned to that response.
@@
    We develop a measurement procedure (RATE) which utilizes two imperfect rewrites to estimate the ATE without needing to explicitly list confounding factors in the evaluation dataset.
@@
    We show that RATE provides a valid estimation of the true causal effect under mild assumptions.
  4. 4.
    We empirically demonstrate that RATE correctly estimates causal effects, while the naive and single-rewrite baselines are highly influenced by spurious associations. We also showt that this causal-vs-correlational problem substantially affects the explanations of leading reward models using common benchmarks.
@@ 2 Setup
A reward function $R$ is a function that takes a prompt $x$ and a response $y$ as inputs and return a real number indicating the quality of the response for the prompt. [^1] A reward function can also be viewed as taking two responses and returning a relative preference between them; our results extend easily to this case as well. To develop an evaluation procedure, we begin with a fixed dataset of prompt-completion pairs $\{(x^{i},y^{i})\}$, where the $x^{i}$ are prompts and the $y^{i}$ are completions (also referred to as “responses”). We are interested in understanding how the reward model responds to a certain attribute $W$ (such as sentiment or length). We consider the case where each prompt-completion pair is labelled with $w^{i}=W(x^{i},y^{i})\in\{0,1\}$ indicating whether the completion has the attribute of interest. [^2] We include the prompt $x$ in the argument of $W$ because the attribute may be prompt-dependent. For example, $W$ might represent helpfulness, which varies based on the context given by the prompt. A recipe could be helpful for questions about cooking but not for questions about history.
We focus on binary attributes for simplicity—many attributes of interest can often be naturally binarized.
Figure 1: Correlations make it hard to isolate the effect of individual attributes on reward models. For instance, if we wish to measure how much a reward model responds to ‘helpfulness’, we must isolate this from correlated attributes (e.g. length).
@@
$\hat{\tau}_{\text{naive}}=\frac{1}{n_{1}}\sum_{\!(x^{i},y^{i}):w^{i}=1}\!R(x^{i},y^{i})-\frac{1}{n_{0}}\sum_{\!(x^{i},y^{i}):w^{i}=0}\!R(x^{i},y^{i})$
@@
This estimates the correlation between the reward and $W$,
$\mathbb{E}[R(X,Y)\mid W=1]-\mathbb{E}[R(X,Y)\mid W=0],$
where the expectation is taken over the distribution from which our evaluation examples are drawn. The problem here is that, even in the infinite data limit, this quantity does not generally isolate the effect of $W$ on $R$. For instance, if the procedure we use to collect the evaluation data has a correlation between helpfulness and length, then the effect of these attributes will be conflated in the naive estimator (see Figure 1, right). Consequently, we will misinterpret the true behavior of the reward model.
@@
$\text{ATE}=\mathbb{E}[R(X,Y(1))-R(X,Y(0))]$ (1)
where $X$ is a random variable for the prompt, and $Y(1)$ and $Y(0)$ are potential outcomes for responses. This quantity is the expected change in reward if we were to change the attribute $W$ from 0 to 1, while keeping all other aspects of the response fixed. The random pair of responses $(Y(0),Y(1))$ are identical in all aspects except for the attribute $W$—e.g., if $W$ is helpfulness then each counterfactual response should have the same writing style, sentiment, topic, etc. In general, we only observe one of the counterfactual responses in our dataset (Figure 1, left)—this is the fundamental problem of causal inference (Imbens & Rubin, 2015).
Choice of Estimand
Beyond the ATE, we will also consider the average treatment effect on the treated (ATT) and the average treatment effect on the untreated (ATU),
ATT $=\mathbb{E}\left[R(X,Y(1))-R(X,Y(0))|W=1\right]$
ATU $=\mathbb{E}\left[R(X,Y(1))-R(X,Y(0))|W=0\right]$
Intuitively, if $W=1$ is a helpful response, the ATT measures the change in reward when we take a helpful response and make it less helpful, and the ATU measures the change in reward when we take an unhelpful response and make it more helpful. These estimands can differ substantially from each other and from the ATE (see Figure 8 in Section D.1). There is no reason to expect these quantities to align in general, so thought should be given to which is most relevant to the question at hand. Indeed, even human preferences are often asymmetric (Kahneman & Tversky, 2013), so we might expect reward model preferences to be as well.
@@
Table 1: Recent language model show promise as (potentially imperfect) rewriters, since qualitatively they are capable of following instructions to change the target attribute (even if they do not always leave off-target attributes unchanged). For instance, here GPT-4o qualitatively does well at rewriting IMDB responses to change sentiment from negative (W = 0) to positive (W = 1). However, this does not ensure that GPT-4o will not change other attributes beside sentiment.
3 Rewrite-based Attribute Treatment Estimator (RATE)
Whatever our choice of estimand, we need a method to estimate it. Here, we develop a method, RATE, that uses rewrites to estimate the causal effect of an attribute on a reward model. The core idea is to create pairs of responses where the only difference is in the attribute of interest, even if the rewrites are imperfect.
@@ 3 Rewrite-based Attribute Treatment Estimator (RATE)
In practice, we implement rewrites using a large language model (LLM) (see Table 1). We begin with a labeled dataset containing ground truth binary variables for attributes such as complexity, sentiment, or helpfulness. We then instruct the LLM to rewrite the responses to the opposite state of the binary variable. For example, we may instruct: “Rewrite this response to express negative sentiment and change nothing else.”
We use $\text{Re}(x^{i},y^{i},w)$ to denote the rewrite operation, which takes a prompt-response pair $(x^{i},y^{i})$ and a desired attribute value $w$, returning a modified response $\tilde{y}^{i}$ such that $W(x^{i},\tilde{y}^{i})=w$. When $x^{i}$ is clear from context, we drop it for brevity as $\text{Re}(y^{i},w)$, even though the rewriter in general may depend on the input prompt.
@@
For instance, when rewriting for “helpfulness”, we might instruct the LLM to “Rewrite this response to be more helpful”, or instruct it to “Rewrite this response to be more helpful, providing additional relevant information or clarification.” In this example, the second instruction makes the meaning of “helpful” more precise. Generally, changing the instruction changes the nature of the rewrites generated, and thus changes the attribute that is being modified.
This is inevitable. Ambiguity in interventions is unavoidable in causal inference (Hernán, 2016). In our context, there is subjectivity in what helpfulness, complexity, or sentiment actually mean. An advantage of rewrite instructions is that we can use natural language to specify, as clearly as possible, what property we are trying to modify. We can understand whether our instructions are having the intended effect by qualitatively examining the rewritten outputs and checking that they indeed change the target attribute. In practice, finding effective rewrite instructions requires an iterative cycle of generating rewrites, examining the responses, and adjusting the rewrite prompt to be more clear and specific.
@@
While qualitative checks help confirm whether the target attribute has changed, they are less effective for detecting off-target modifications—unintended edits to grammar, tone, or other attributes (see Table 2).
For example, in Table 1 (and Section E.3), sentiment flips successfully, but other properties may also shift. This issue can be identified quantitatively by comparing reward distributions before and after rewriting (Figure 2). Ideally, if only sentiment changed, a second rewrite reversing the attribute should restore the original distribution. Otherwise, the (original, rewrite) pairs are not perfectly counterfactual and cannot directly estimate Equation 1.
Mathematically, rewriting introduces some error $\epsilon^{i}_{w}$ in the observed reward because we cannot perfectly construct the true counterfactual $y^{i}(w)$, which should only differ in the target attribute:
$\epsilon^{i}_{w}=R(x^{i},\text{Re}(y^{i},w))-R(x^{i},y^{i}(w))$
We would like to correct for these errors. Yet, the whole point of the rewrites is to approximate the counterfactuals $y^{i}(w)$, so we cannot directly measure $\epsilon^{i}_{w}$.
Figure 2: Off-target changes from imperfect rewrites affect the reward measurement. Ideally, if rewrites affected only the target attribute (sentiment), then applying a second rewrite to revert the change should restore the original reward distribution. Unfortunately, the observed distribution shift indicates that off-target modifications occur during rewriting. Here, the original samples (blue) are drawn from the HH-RLHF dataset, and are rewritten twice on sentiment (orange). Rewards are from ArmoRM.
@@
Our solution is to introduce more noise. Instead of comparing a rewrite to the original response, we compare it to the rewrite of the rewrite, thereby canceling out off-target noise introduced by the rewrite process. That is, rather than selecting (original, rewrite):
$\tilde{\delta}^{i}=\begin{cases}R(x^{i},y^{i})-R(x^{i},\text{Re}(y^{i},0)),&\text{if }w^{i}=1\\ R(x^{i},\text{Re}(y^{i},1))-R(x^{i},y^{i}),&\text{if }w^{i}=0\end{cases}$
@@
$\delta^{i}=\begin{cases}&R(x^{i},\text{Re}(\text{Re}(y^{i},0),1))-R(x^{i},\text{Re}(y^{i},0)),\\ &\quad\quad\text{if }w^{i}=1\\[4.30554pt] &R(x^{i},\text{Re}(y^{i},1))-R(x^{i},\text{Re}(\text{Re}(y^{i},1),0)),\\ &\quad\quad\text{if }w^{i}=0\end{cases}$
The motivation is that the off-target changes introduced by the rewrite process will, in expectation, cancel out when we are comparing two things in “rewrite space”. For example, the tendency for LLMs to produce well-formatted text will affect both the first rewrite and the rewrite of the rewrite (see Table 2), so the overall contribution of this off-target change will cancel out.
In Algorithm 1 we use this idea to define the Rewrite-based Attribute Treatment Estimators (RATE) for the ATT, ATU, and ATE. These estimators are simply the averages of the reward difference between the rewrites and the rewrites of the rewrites.
@@
1:  Input: Dataset $\{(x^{i},y^{i},w^{i})\}$, reward model $R$, function $\text{Re}()$
2:  Return: Estimates $\widehat{\text{ATT}}_{\text{RATE}}$, $\widehat{\text{ATU}}_{\text{RATE}}$, $\widehat{\text{ATE}}_{\text{RATE}}$
3:  Initialize $n_{1}\leftarrow\sum_{i,j}\mathbb{I}[w^{i}=1]$, $n_{0}\leftarrow\sum_{i,j}\mathbb{I}[w^{i}=0]$
4:  $\widehat{\text{ATT}}_{\text{RATE}}\leftarrow\frac{1}{n_{1}}\sum\limits_{i:w^{i}=1}[R(x^{i},\text{Re}(\text{Re}(y^{i},0),1))-R(x^{i},\text{Re}(y^{i},0))]$
5:  $\widehat{\text{ATU}}_{\text{RATE}}\leftarrow\frac{1}{n_{0}}\sum\limits_{i:w^{i}=0}[R(x^{i},\text{Re}(y^{i},1))-R(x^{i},\text{Re}(\text{Re}(y^{i},1),0))]$
6:  $\widehat{\text{ATE}}_{\text{RATE}}\leftarrow\frac{n_{1}}{n_{0}+n_{1}}\widehat{\text{ATT}}_{\text{RATE}}+\frac{n_{0}}{n_{0}+n_{1}}\widehat{\text{ATU}}_{\text{RATE}}$
7:  Return: $\widehat{\text{ATT}}_{\text{RATE}}$, $\widehat{\text{ATU}}_{\text{RATE}}$, $\widehat{\text{ATE}}_{\text{RATE}}$
4 Cancellation of Multiple Rewrite Errors
We now turn to the validity of the RATE procedure. Intuitively, the idea is that as long as the errors introduced by the rewrite are independent of the target attribute $W$, then these errors will cancel out when we compute the average. A particularly important special case of independence is when the LLM always introduces the same error, irrespective the rewrite direction. For example, we observe that GPT-4o rewrites nearly always correct typos, independent of the attribute being rewritten. In this case, the single-rewrite estimator is biased because it includes the effect of the typo correction, but RATE is not because it compares typo-corrected responses to typo-corrected responses. We now formalize these intuitions.
Latent Variable Model for the Responses
We can partition all the possible high-level attributes of a response into three categories: the target attribute $W$ we want to change (e.g. sentiment), off-target attributes $Z$ which are always unaffected by the rewrite (e.g. topic, and language), and off-target attributes which might end up being affected by the rewrite (e.g. grammatical structure), which we’ll denote $\xi$. We write the response as $Y=Y(W,Z,\xi)$.
Assumption 1: Rewrite Errors Don’t Noticeably Depend on the Rewrite Direction
The off-target changes, $\tilde{\xi}$, introduced by the rewrite process are randomly drawn from some distribution, $P_{\text{Re}}$ which doesn’t depend on $W$:
[…]
with $\text{with }\tilde{\xi}$ $\sim P_{\text{Re}}(\tilde{\xi})$
Continuing the example where rewriting sentiment ($W$) corrects typos ($\xi$) but leaves language ($Z$) unchanged, this states that GPT-4o isn’t more or less likely to fix typos on positive sentiment responses compared to negative sentiment. Note that this is in fact weaker than assuming that LLM-rewriters tend to repeat the same rewrite errors (e.g. GPT-4o always fixes typos no matter what). If we do observe this behavior, then the assumption is trivially satisfied.
Assumption 2: Additivity of Reward with Respect to Rewrite Errors
Suppose that the reward function can be decomposed additively into two components: